OpenStudy (anonymous):

A= ? b=24 c=40? Use a squared + b squared= c squared to find a

OpenStudy (anonymous):

that means apply the pyth theorem a^2 + b^2 = c^2 a^2 = c^2 - b^2 a = sqrt (c^2 - b^2) plug and calculate